A company manufactures in a foreign country and sells only to US customers so that revenue is in USD and variable costs are in foreign currency (FC). The company anticipates the following results in the coming year:

  • Sales of 7,500 units

  • Price of $250 per unit

  • Per unit cost of FC310 per unit

  • Exchange rate of 1.50 FC/$

  • Additional fixed costs of $150,000

  • No interest expense (100% equity capital structure)

  • Income tax rate of 25%

    Build an income statement for the company using these figures.

  • What are the anticipated net income and profit margin for the coming year?

  • What is the breakeven number of units sold?
Answer #1

Preparing the Income Statement of the company for the coming year;-

Particular Amount in $
Sales (7500*250) 1875,000
Less: Cost [(FC310/1.50 FC per $)*7500] (1550,000)
Gross Income 325,000
Less: Additional Fixed Cost (150,000)
Operating Income 175,000
Less: Tax@25% (43,750)
Net Income 131,250

Anticipated Net Income for the coming year = $ 131,250

Anticipated Profit margin for the coming year= (Net Income/Net Sales)*100

= (131,250/1875,000)*100

= 7%

Break-even number of units sold = Fixed Cost/(Sales per unit- Variable cost per unit)

= 150000/(250-206.67*)

= 3461.80 or 3462 units

*- Variavle Cost per unit = FC310/1.50 FC per $ ;= $206.67

If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating.